OpenStudy (anonymous):

Which functions below have period ? y = cos x y = sin x y = sec x y = tan x y = cot x y = csc x

geerky42 (geerky42):

All of them do. Not all have the same period, though.

OpenStudy (anonymous):

the period 2 pi. sorry if you could please explain how to find the period I would greatly appreciate it

jimthompson5910 (jim_thompson5910):

for sine and cosine, the general form is y = A*sin(Bx-C) + D y = A*cos(Bx-C) + D ------------------------------------------------------- in both cases, the period is T = 2pi/B

jimthompson5910 (jim_thompson5910):

so for instance in y = 2*sin(3x-7) B = 3 Period T = 2pi/B T = 2pi/3